The length of a rectangle is twice the width. The perimeter of the rectangle is 41.1 . What is the width?
Answer by addingup(3677)

L = 2W
2L+2W = 41.1
2(2W)+2w = 41.1
4W+2W = 41.1
6W = 41.1
W = 6.85 and Length:
L = 2*6.85 = 13.7
----------------------
Check:
2(13.7)+2(6.85)
= 27.4+13.70 = 41.1 Correct
:
John